How many moles of O2 are required to react with 2 moles of H2?


1
Expert's answer
2021-02-02T01:21:26-0500

Solution:

Typically H2 and O2 react to form water (H2O).

The balanced chemical equation:

2H2 + O2 → 2H2O

According to the equation above: n(H2)/2 = n(O2)

Hence,

n(O2) = n(H2)/2 = 2 mol / 2 = 1 mol

n(O2) = 1 mol


Answer: 1 mol of O2 are required.